OpenStudy (anonymous):

sketch graph of function f(x)=x^2/(x^2-4) using any extrema, intercepts, symmetry, and asymptotes

OpenStudy (danjs):

ok, well you are going to need the first and second derivatives , so why not start out by finding those

OpenStudy (anonymous):

yeah i know, currently finding both the derivatives

OpenStudy (danjs):

\[\frac{ d }{ dx }\frac{ x^2 }{ (x^2-4) }\]

OpenStudy (anonymous):

\[\frac{ -8 }{ (x ^{2}-4 )^{2}}\]

OpenStudy (anonymous):

*-8x

OpenStudy (danjs):

From the denominator of f(x), it factors to (x+2)(x-2), so if the denominator is zero, the function does not exist, for X=2 and X=-2 you have vertical asymptotes

OpenStudy (danjs):

\[\lim_{x \rightarrow +- \infty} f(x) = \]

OpenStudy (danjs):

As x goes off to infinity or negative infinity, the function f(x) goes to 1 Horizontal asmytope at Y=1

OpenStudy (danjs):

\[\lim_{x \rightarrow \infty} \frac{ 1 }{ 1 - \frac{ 4 }{ x^2 } }\]

OpenStudy (danjs):

divided everything by x^2 , then let x go to infinity,

OpenStudy (anonymous):

oh k, i already know of the asmytopes, jut having trouble on finding the derivatives, and the intercepts

OpenStudy (danjs):

ok...

OpenStudy (anonymous):

mostly the second deriv

OpenStudy (danjs):

The first derivative is missing a x term in the numerator, \[y ' = \frac{ -8x }{ (x^2-4)^2 }\]

OpenStudy (danjs):

using the quotient rule

OpenStudy (danjs):

oh i see you corrected it

OpenStudy (danjs):

so \[y '' = \frac{ d }{ dx }\frac{ -8x }{ (x^2-4)^2 }\]

OpenStudy (danjs):

same quotient rule as the first derivative gives... \[y '' = \frac{ 8(3x^2 +4) }{ (x^2 - 4)^2 }\]

OpenStudy (anonymous):

yeah i keep ending up a ridiculous number as an answer for the second derivative, such that i'm struggling on it

OpenStudy (anonymous):

oh k thank you

OpenStudy (danjs):

you could let u= x^2 + 4 du = 2x dx

OpenStudy (anonymous):

oh yeah, i forgot that i could do that huh

OpenStudy (danjs):

ok so now you need to test for Increasing/Decreasing Concave up/Concave Down

OpenStudy (anonymous):

i know that the first one is the critical of 0, 2

OpenStudy (danjs):

well 2 is a vertical asmtope

OpenStudy (danjs):



OpenStudy (danjs):

That is what we have so far, with x=0 a critical point

OpenStudy (anonymous):

oh k, then wouldn't there be zero inflection?